. The radioactive isotope carbon-10 has a half-life of 20 seconds.

a. How much time is required so that only 1/16 of the original amount remains?

b. Find the rate of decay at this time.


1
Expert's answer
2022-04-11T18:15:39-0400

"T_{1\/2}=20secs"

"\lambda=ln2/T_{1/2}"

a. "1 \rightarrow 1/2" takes 20seconds

"1/2 \rightarrow 1/4" takes 20 seconds

"1/4 \rightarrow 1/8" takes 20 seconds

"1/8 \rightarrow 1/16" takes 20 seconds

It takes 80secs for 1/16of the original to remain.

b."decay\ rate=-\frac{dN}{dt}=\lambda N"

10g=6.03×1023 molecules

decay rate=0.035×6.03×10^{23}=2.1×10^{22}0.035×6.03×1023 =2.1×1022

Particles per sec